OpenStudy (anonymous):

how do you solve (-3-4x)^1/2-(-2-2x)^1/2=1

OpenStudy (anonymous):

alright, if I'm not wrong , you'll have to multiply the left side with \[\sqrt{-3-4x} +\sqrt{-2-2x}\] up and down , then you'll get : \[(-3x-4x) -(-2-2x) = 1\] simply subtract and solve for x and you'll get : \[-3-4x +2 +2x -1 = 0\] -2x -1 -1 = 0 -2x = 2 x = -1 Correct me if I'm wrong ^_^

OpenStudy (anonymous):

oh thank you so much thank you =) i was really having a hard time with this

OpenStudy (anonymous):

and i will

OpenStudy (anonymous):

yeah i think your right

OpenStudy (anonymous):

you're welcome ^_^ good luck!